# System for Systematic Literature Review Using Multiple AI agents: Concept and an Empirical Evaluation

### v1

@@ 8 Conclusions
The development and implementation of our multi-AI agent model represent a significant advancement in the field of SLR. By integrating the capabilities of LLMs, this research demonstrates a novel approach to automating and optimizing the SLR process. Our model addresses the primary challenges associated with traditional SLR methods: the time-consuming nature of the process and the potential for human error or bias in literature selection and analysis. By automating the initial search, screening, summarization, and analysis phases, the model significantly reduces the manual effort and time required, while also enhancing the accuracy and consistency of the results.
The use of a simple user interface for topic input and subsequent generation of tailored search strings illustrates the model’s user-friendly approach, making complex SLR processes accessible to a broader range of researchers. The inclusive and exclusive filtering mechanism ensures that the literature review remains focused and relevant, directly aligning with the specified research questions. The autonomous summarization of abstracts and the final analytical phase underscore the model’s ability to extract vast amounts of data into clear, relevant information, a task that would be challenging without the aid of advanced AI.
This research contributes to the growing field of AI application in academic research, showcasing how LLMs can be effectively employed to enhance research methodologies. While the model significantly improves efficiency and accuracy, it is important to acknowledge the role of human oversight in guiding and interpreting the results, ensuring that the final output maintains the depth required in scholarly research.

### v2

@@ 8 Conclusions
The development and implementation of our multi-AI agent system represent a significant advancement in the field of SLR. By integrating the capabilities of LLMs, this research demonstrates a novel approach to automating and optimizing the SLR process. Our system addresses the primary challenges associated with traditional SLR methods: the time-consuming nature of the process and the potential for human error or bias in literature selection and analysis. By automating the initial search, screening, summarization, and analysis phases, the system significantly reduces the manual effort and time required, while also enhancing the accuracy and consistency of the results.
The use of a simple user interface for topic input and subsequent generation of tailored search strings illustrates the system’s user-friendly approach, making complex SLR processes accessible to a broader range of researchers. The inclusive and exclusive filtering mechanism ensures that the literature review remains focused and relevant, directly aligning with the specified research questions. The autonomous summarization of abstracts and the final analytical phase underscore the system’s ability to extract vast amounts of data into clear, relevant information, a task that would be challenging without the aid of advanced AI.
This research contributes to the growing field of AI application in academic research, showcasing how LLMs can be effectively employed to enhance research methodologies. While the system significantly improves efficiency and accuracy, it is important to acknowledge the role of human oversight in guiding and interpreting the results, ensuring that the final output maintains the depth required in scholarly research.